OpenStudy (anonymous):

HELP PLEASE<3 WILL GIVE MEDAL!(: Task 1: Solve using 2 different methods. Explain which method you found to be more efficient. a. 3x - 9y = 3 6x - 3y = -24

OpenStudy (anonymous):

b. 7x - 3y = 20 5x + 3y = 16 c. y = 1/2x - 6 2x + 6y = 19

OpenStudy (anonymous):

what grade?

OpenStudy (anonymous):

9th , I think. i'm in 10th but, I have to make up my credits from freshman year